Operations Report – May 14th

Crew 212 Operations Report 14-MAY-2019

SOL: 9

Name of person filing report: Paolo Guardabasso

Non-nominal systems: Rover Opportunity.

Notes on non-nominal systems: the rover Opportunity failed again this morning, during the engineering EVA that was performed to test the three rovers (see the dedicated EVA report).

Generator (hours run last night): 12:00

On (last night): 7:00 PM

Off (this morning): 7:00 AM

Solar— SOC 59% at 7:00 PM

Diesel Reading – 88%

Propane Reading – 46%

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 300 gallons (the value of yesterday was 380, it was probably overestimated, as the water meter says we consumed 41 gallons today)

Axillary to Static tank transfer – no

Gallons transferred: No

Water in GreenHab – 130 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 143624.1

Toilet tank emptied: Yes

Deimos rover used: no (out of order)

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes

Hours: 93.3h

Beginning charge: 100%

Ending charge: 65%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 59.8h

Beginning charge: 100%

Ending charge: 99%

Currently charging: No

Curiosity rover used: Yes

Hours: 96.6h

Beginning charge: 100%

Ending charge: 87%

Currently charging: Yes

Notes on rovers: a 4 miles rover test was run during an Engineering EVA (a dedicated report will be sent to Mission Support); Opportunity stopped after 1.3 miles for battery overheating (similar to Sol 8).

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? Nothing to report.

CrewCar used and why, where? Off Campus

General notes and comments: the crew of EVA#10 had a delay on their scheduled return time. We received the new loft pump and we’ll assemble it later tonight or tomorrow morning.

Summary of internet: Nothing to report.

Summary of suits and radios: radio batteries are often giving wrong values (indicating full battery and then turning off). From now on, spare batteries will be brought during EVAs. See spacesuit report for more info.

Summary of Hab operations: The ceiling in the commander’s room seems dryer than yesterday night, thanks to the temporary repair done yesterday.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: We’re about to run out of milk (2 small bags), carbs (no pasta, 1 portion each of rice), vegetables (no broccoli, only corn and onion left).